# Objective Evaluation of Prosody and Intelligibility in Speech Synthesis via Conditional Prediction of Discrete Tokens

### v1

@@ 5.C DISCUSSION ON INTELLIGIBILITY & QUALITY
Table 2: Correlations of objective metrics with MOS in SOMOS and VoiceMOS datasets.
Metric Speech token SOMOS VoiceMOS22
Utt-level Sys-level Utt-level Sys-level
LCC SRCC LCC SRCC LCC SRCC LCC SRCC
WER Word 0.32 0.30 0.62 0.63 0.30 0.31 0.42 0.41
CER Character 0.34 0.33 0.60 0.68 0.28 0.31 0.40 0.48
SpeechLMScore [19]
HuBERT
(k=50)
L3 0.04 0.04 0.12 0.08 0.48 0.50 0.56 0.65
uLM(ours)
HuBERT
(k=50)
L3 0.06 0.06 0.17 0.14 0.46 0.46 0.53 0.63
uLM (ours)
HuBERT
(k=500)
L3 0.17 0.16 0.37 0.36 0.57 0.57 0.69 0.72
L9 0.34 0.33 0.62 0.62 0.47 0.48 0.61 0.63
L12 0.24 0.22 0.50 0.50 0.44 0.44 0.54 0.56
TTScore-int (proposed)
HuBERT
(k=50)
L3 0.37 0.35 0.64 0.60 0.42 0.47 0.48 0.56
TTScore-int (proposed)
HuBERT
(k=500)
L3 0.40 0.38 0.69 0.68 0.48 0.52 0.54 0.64
L9 0.43 0.42 0.72 0.71 0.42 0.46 0.50 0.58
L12 0.42 0.40 0.71 0.70 0.41 0.44 0.48 0.56
@@ 6 RESULTS FOR PROSODY EVALUATION
(a) Score distributions for real vs synthesized speech
(b) Score distributions with original vs perturbed pitch (HifiTTS)
@@ 6.A REAL VS SYNTHESIZED SPEECH SCORES
Table 3: Prosody correlations across benchmarks
@@
Log f0 RMSE -0.03 -0.03 -0.08 -0.09
f0-corr 0.02 0.04 0.07 0.02
TTScore-pro (small) 0.05 0.04 0.23 0.15
TTScore-pro (large) 0.06 0.05 0.25 0.18
@@
All Systems 0.33 0.33 0.45 0.46
Top 25% Systems 0.20 0.21 0.25 0.32
Top 12% Systems 0.19 0.19 0.30 0.24
(c) Correlation with TTSArena ELO System-level Metric LCC SRCC Log f0 RMSE 0.18* 0.32* f0-corr 0.13 0.32 TTScore-pro (large) 0.25 0.21
@@ 6.B ORIGINAL VS PERTURBED PITCH SCORES
The score distributions for resynthesized speech with correct and perturbed prosody in Figure 2(b) and 2(c) show that resynthesized speech with original prosody has higher scores than resynthesized speech with perturbed pitch, consistently in both scenarios. Since pitch contours being disturbed or not is the only varying factor between the cases, the score distribution analysis proves that the prosody metric accounts for prosody appropriateness, showing lower values for disrupted pitch.
@@ 6.C CORRELATION WITH MOS AND TTSArena ELO
Since lower F0-RMSE indicates better prosody, a negative correlation with MOS is expected, while positive correlations are expected otherwise. As both baselines require reference speech, results are reported only on SOMOS, while VoiceMOS lacks aligned references. As shown in Table 3(a), the baselines show weak or no correlation with MOS, confirming their limited reliability for reflecting perceived naturalness.
In contrast, TTScore-pro achieves consistently stronger correlations. While moderate in absolute terms, this is expected since naturalness depends on multiple factors beyond prosody. Results from VoiceMOS (Table 3(b)) further confirm its robustness, likely due to the dataset’s greater diversity in synthesis methods and prosody modeling quality. Overall, TTScore-pro provides a more reliable and informative objective prosody metric than traditional measures.
Table 3(c) shows correlations with TTSArena ELO scores, where higher ELO indicates better perceived quality. While F0-RMSE should correlate negatively with ELO, it unexpectedly shows a positive trend, confirming its unreliability as a prosody metric. F0-corr performs slightly better but remains inconsistent across datasets. In contrast, TTScore-pro consistently aligns with human preferences in this benchmark, showing that it captures prosodic naturalness more reliably. These results highlight TTScore-pro as a more robust and informative objective metric, compared to traditional pitch-based measures which fail to reflect overall perception.
@@ 7 CONCLUSION
In this work, we introduced a targeted evaluation framework for synthesized speech that measures intelligibility and prosody as distinct aspects of speech quality. We proposed two conditional likelihood–based metrics: TTScore-int, which evaluates how well the linguistic content of speech is preserved, and TTScore-pro, which assesses the appropriateness of prosody relative to the intended content. By leveraging specialized discrete speech tokens and text-conditioned sequence-to-sequence predictors, both metrics provide reference-free, speech-domain evaluations that overcome many of the shortcomings of conventional approaches such as WER, CER, and F0-based measures.
Comprehensive experiments across SOMOS, VoiceMOS, and TTSArena benchmarks demonstrated that TTScore-int and TTScore-pro not only correlate strongly with established baselines but also align more closely with human judgments of naturalness. These results confirm the value of targeted evaluation: aspect-specific metrics provide finer insights while remaining predictive of overall perceived quality. This work opens new opportunities for extending targeted evaluation beyond intelligibility and prosody. The proposed paradign of conditional generation for evaluation is highly flexible for new formulations for different aspects of speech.

### v2

@@ 5.B CORRELATION WITH MOS
Additionally, we include UTMOS [34], the top-performing system from the VoiceMOS 2022 Challenge, as a reference. Results are reported on the VoiceMOS test subset, as UTMOS is trained on the VoiceMOS 22 training data. As expected, UTMOS achieves high correlation with MOS on VoiceMOS, but its performance is notably worse on the SOMOS dataset. Although UTMOS is trained with MOS-labeled data and directly targets MOS prediction, it shows lower correlation with MOS than the proposed metric on SOMOS, highlighting the challenge of cross-dataset generalization for MOS predictors that rely on MOS-labeled data, which is often limited in scale. In contrast, the proposed aspect-specific evaluation is not limited by the need for MOS-labeled data and exhibits more stable behavior across datasets, likely due to its more focused and simpler objective.
(a) Score distributions for real vs synthesized speech
(b) Score distributions with original vs perturbed pitch (HifiTTS)
@@ 6.A REAL VS SYNTHESIZED SPEECH SCORES
Table 3: Prosody correlations across benchmarks (with bootstrapped 95% confidence intervals)
@@
Log f0 RMSE -0.03 (-0.09, 0.03) -0.03 (-0.09, 0.03) -0.08 (-0.22, 0.06) -0.09 (-0.23, 0.05)
f0-corr 0.02 (-0.04, 0.08) 0.04 (-0.02, 0.10) 0.07 (-0.07, 0.21) 0.02 (-0.11, 0.17)
TTScore-pro (small) 0.05 (0.00, 0.11) 0.04 (-0.02, 0.10) 0.23 (0.10, 0.35) 0.15 (0.03, 0.26)
TTScore-pro (large) 0.06 (0.00, 0.12) 0.05 (-0.01, 0.11) 0.25 (0.11, 0.38) 0.18 (0.04, 0.32)
@@
All Systems 0.33 (0.31, 0.35) 0.33 (0.31, 0.36) 0.45 (0.32, 0.57) 0.46 (0.31, 0.60)
Top 25% Systems 0.20 (0.14, 0.26) 0.21 (0.15, 0.27) 0.25 (0.13, 0.34) 0.32 (0.21, 0.41)
Top 12% Systems 0.19 (0.10, 0.28) 0.19 (0.10, 0.28) 0.30 (0.15, 0.44) 0.24 (0.10, 0.39)
(c) Correlation with TTSArena ELO
System-level
Metric LCC SRCC
Log f0 RMSE 0.18* (-0.51, 0.73) 0.32* (-0.40, 0.86)
f0-corr 0.13 (-0.55, 0.70) 0.32 (-0.36, 0.89)
TTScore-pro (large) 0.25 (-0.45, 0.76) 0.21 (-0.50, 0.73)
@@ 6.B ORIGINAL VS PERTURBED PITCH SCORES
The score distributions for resynthesized speech with correct and perturbed prosody in Figure 2(b) and 2(c) show that resynthesized speech with original prosody has higher scores than resynthesized speech with perturbed pitch, consistently in both scenarios. Since pitch contours being disturbed or not is the only varying factor between the cases, the score distribution analysis proves that the proposed prosody metric accounts for pitch appropriateness, showing lower values for disrupted pitch.
@@ 6.C CORRELATION WITH MOS AND TTSArena ELO
Since lower F0-RMSE indicates better pitch-related prosody, a negative correlation with MOS is expected, while positive correlations are expected otherwise. As both baselines require reference speech, results are reported only on SOMOS, while VoiceMOS lacks aligned references. As shown in Table 3(a), the baselines show weak or no correlation with MOS, confirming their limited reliability for reflecting perceived naturalness.
In contrast, TTScore-pro achieves consistently stronger correlations. While moderate in absolute terms, this is expected since naturalness depends on multiple factors beyond prosody. The correlations are lower in general compared to TTScore-int, which can be attributed to intelligibility’s more obvious effect on overall perception and higher complexity of prosody modeling. Results from VoiceMOS (Table 3(b)) further confirm its robustness, likely due to the dataset’s greater diversity in synthesis methods and prosody modeling quality. Overall, TTScore-pro provides a more reliable and informative objective prosody metric than traditional measures.
Table 3(c) shows correlations with TTSArena ELO scores, where higher ELO indicates better perceived quality. While F0-RMSE should correlate negatively with ELO, it unexpectedly shows a positive trend, confirming its unreliability as a prosody metric. F0-corr performs slightly better but remains inconsistent across datasets. In contrast, TTScore-pro consistently aligns with human preferences in this additional benchmark, the results indicate that it tends to capture the appropriateness of pitch patterns and tend to reflect prosodic naturalness more reliably. We note that this benchmark has fewer number of synthesis systems. Overall results highlight TTScore-pro as a more robust and informative objective metric, compared to traditional pitch-based measures which fail to reflect overall perception.
7 LIMITATIONS
While we examine a very important aspect of prosody, namely F0, prosody is a broad concept that also encompasses other speech attributes such as rhythm, energy, and phrasing. In this work, we focus on F0 due to its importance and because it is the most commonly measured prosodic feature in speech synthesis.
Moreover, since token distributions are learned from data, our method is data-driven and may be susceptible to underlying biases. In this work, we utilize a large-scale English dataset recorded mostly under clean and controlled conditions. We focus on this setting, as such conditions are most widely used and commonly focused on in speech synthesis research and benchmarking. However, our method may have limitations in scenarios where recording conditions are less controlled (e.g., spontaneous speech or in-the-wild settings) and does not currently support other languages. We also note that, since our method does not require labels beyond text transcriptions, which also can be obtained automatically, it is directly applicable to other scenarios where such data is available.
@@ 8 CONCLUSION
In this work, we introduced a targeted evaluation framework for synthesized speech that measures intelligibility and prosody as distinct aspects of speech quality. We proposed two conditional likelihood–based metrics: TTScore-int, which evaluates how well the linguistic content of speech is preserved, and TTScore-pro, which assesses the appropriateness of pitch, a crucial aspect of prosody, relative to the intended content. By leveraging specialized discrete speech tokens and text-conditioned sequence-to-sequence predictors, both metrics provide reference-free, speech-domain evaluations that overcome many of the shortcomings of conventional approaches such as WER, CER, and F0-based measures.
Comprehensive experiments across SOMOS, VoiceMOS, and TTSArena benchmarks demonstrated that TTScore-int and TTScore-pro not only correlate strongly with established baselines but also align more closely with human judgments of naturalness. These results confirm the value of targeted evaluation: aspect-specific metrics provide finer insights while remaining predictive of overall perceived quality. This work opens new opportunities for extending targeted evaluation beyond intelligibility and prosody. The proposed paradigm of conditional generation for evaluation is highly flexible for new formulations for different aspects of speech.